OpenStudy (anonymous):

Let F(x)=x^2-1 and G(x)=3-x. Find (F-G)(6)

OpenStudy (anonymous):

Let F(x)=x^2-1 and G(x)=3-x. Find (F-G)(6) \[F(x)=x^2-1\] \[G(x)=3-x\] \[F(x)-G(x)=(F-G((x)=x^2-1 -(3-x) =x^2 + x -2 \] \[(F-G))(6)= =6^2 + 6 -2 =40 \]

OpenStudy (anonymous):

\[ F(x)-G(x)=(F-G((x)=x^2-1 -(3-x) =x^2 -1-3+x=x^2+x-4 \]

OpenStudy (anonymous):

\[ (F-G))(6)= =6^2 + 6 -4 =38 \]

OpenStudy (anonymous):

@abtster

OpenStudy (anonymous):

Yes you're right, me bad

OpenStudy (anonymous):

Ohh this arithmetic of mine